# Evaluating the Robustness of Adversarial Defenses in Malware Detection Systems

### v1

@@ Abstract
Machine learning (ML) has become a vital tool in Android malware detection, effectively identifying malicious patterns in applications to combat the growing threat of malware. However, ML-based systems remain highly vulnerable to evasion attacks, where subtle perturbations can bypass detection mechanisms. Despite efforts to develop adversarially robust malware detection systems, the lack of comprehensive evaluation frameworks for state-of-the-art defenses in this domain limits the thorough understanding of their effectiveness. This study introduces two key contributions to advance adversarial evaluation in binary-constrained domains. First, we propose Prioritized Binary Rounding, a technique that efficiently transforms continuous adversarial perturbations into binary spaces while preserving high attack success rates and minimizing perturbation size. Second, we develop the $\sigma$-binary attack, a novel adversarial attack specifically tailored for binary domains, designed to achieve adversarial objectives with minimal perturbations while maintaining compatibility with binary feature constraints. Extensive evaluations on the Malscan dataset demonstrate the superiority of the proposed approach. $\sigma$-binary consistently outperforms existing adversarial methods, including CW, $\sigma$-zero, Mimicry, and PGD-based attacks, across both robust and non-robust defenses in terms of attack success rate and perturbation size. Furthermore, our analysis of state-of-the-art defenses using $\sigma$-binary reveals critical vulnerabilities. Defenses equipped with adversary detectors, such as KDE, DLA, DNN+, and ICNN, exhibit significant brittleness, with attack success rates exceeding 90% for fewer than 10 feature modifications and reaching 100% with a 20-feature perturbation budget. Adversarially trained defenses, including AT-rFGSMk and AT-MaxMA, exhibit improved robustness at lower budgets but remain highly susceptible to unrestricted perturbations, with attack success rates of 99.45% and 96.62%, respectively. Although PAD-SMA has been demonstrated to maintain a high accuracy above 83.45% (equivalent to an attack success rate below 16.55%) against 15 gradient-based attacks without perturbation limits, our analysis reveals significant vulnerabilities. The $\sigma$-binary attack achieves a 36.34% success rate against PAD-SMA with fewer than 10 modified features, increasing to 94.56% under unrestricted perturbations. These findings underscore the importance of precise evaluation tools, such as $\sigma$-binary, to uncover latent weaknesses in existing defenses and guide the development of more robust malware detection systems.
Index Terms:
Machine Learning, Android Malware Detection, Adversarial Malware Detection, Evasion Attack
@@ I Introduction
The surge in mobile device attacks has escalated into a critical global concern. In 2023 alone, such attacks increased by 50% compared to the previous year. According to Kaspersky, their systems blocked 33.8 million instances of malware, adware, and riskware targeting mobile devices, with Android remaining the primary focus of cybercriminals [1]. This significant increase in Android malware underscores the urgent need for advanced detection strategies to counter increasingly sophisticated threats.
@@
Machine learning (ML) has become a cornerstone of Android malware detection, significantly improving scalability and detection precision [2, 3, 4, 5, 6]. Among ML-based approaches, static analysis methods, such as Drebin [4] and MaMaDroid [7], have demonstrated notable efficiency. These methods analyze application features without requiring execution, enabling rapid and scalable security assessments. Consequently, static analysis has become a foundational component of modern malware detection pipelines.
@@
Despite these advancements, ML-based malware detectors are inherently vulnerable to adversarial attacks. These attacks exploit weaknesses in detection models by introducing carefully crafted perturbations to evade detection. Such attacks, which undermine reliability, are broadly categorized into evasion and poisoning attacks. Evasion attacks manipulate test examples during the deployment phase to bypass detection [8, 9, 10, 11, 12], whereas poisoning attacks compromise training datasets to degrade detector performance [13, 14]. This study focuses on evasion attacks because of their immediate implications for operational systems.
@@
Evasion attacks can be further divided into problem-space and feature-space attacks. Problem-space attacks modify malware applications directly, such as injecting benign ”gadgets” or altering application components to obscure malicious intent [15, 16, 17]. In contrast, feature-space attacks modify feature vectors extracted from malware while preserving the functionality of the original sample [9, 8]. This study focuses exclusively on feature-space attacks because of their controlled evaluation capabilities and practical relevance.
@@
Defending against adversarial attacks remains a significant challenge. Adversarial training, which incorporates adversarial samples into training datasets, has shown promise in improving robustness [8]. However, these methods often fail to generalize to novel attack strategies, impose high computational costs, and risk overfitting to specific attacks [18]. Other defense mechanisms such as ensemble methods [19, 20, 21], auxiliary models such as Variational Autoencoders (VAEs) [22], and hybrid frameworks such as PAD-SMA [23] offer varying degrees of effectiveness. However, these approaches frequently involve trade-offs, including increased complexity and reduced detection accuracy.
I-A Our Contributions
@@
Evaluating the adversarial robustness of malware defenses in binary space presents significant challenges. Existing approaches encounter two critical limitations. First, gradient-based attacks originally designed for continuous spaces are often adapted to binary domains using naive binarization techniques such as Deterministic and Randomized Binary Rounding, which significantly reduce their effectiveness (e.g., [9]). Additionally, the absence of a gradient-based attack designed for binary spaces arises from the inherent complexity of optimization within a non-convex and non-differentiable constraint space, further exacerbating this limitation. Second, these methods rely on subjective methodologies and generalized attack strategies, employing low-iteration gradient-based attacks with fixed hyperparameters across all defenses without optimizing them to individual models (e.g., [23, 22, 24, 20, 25]). This combination of attack incompatibility with binary-constrained domains and insufficient optimization prevents these methods from effectively exposing vulnerabilities in defense mechanisms.
To address the challenges of evaluating adversarial robustness in binary-constrained domains, this study introduces a novel framework tailored for thorough assessment of defenses. The key contributions of our research are as follows:
First, we propose Prioritized Binary Rounding, an innovative technique for transforming gradient-based attacks from continuous to binary spaces. Extensive evaluations on the Malscan dataset [26], encompassing four defenses (DNN, AT-rFGSM, AT-MaxMA, and PAD-SMA) and two attack scenarios ($\sigma$-zero and CW attacks, optimizing $\ell_{0}$-norm and $\ell_{2}$-norm, respectively), demonstrate its superiority over Deterministic and Randomized Binary Rounding in terms of both attack success rate and perturbation size IV-B. For the $\sigma$-zero attack, Prioritized Binary Rounding achieves results nearly identical to continuous outputs, minimizing feature modifications while maintaining high attack success rates: $100\%$ on DNN (compared with 94.88% for the second-best method), $85.95\%$ on AT-rFGSM (vs. $78.64\%$), $92.56\%$ on AT-MaxMA (vs. $81.31\%$), and $79.64\%$ on PAD-SMA (vs. $69.91\%$). For the CW attack, it effectively addresses the challenge of binarizing small fractional perturbations, achieving $100\%$ success on DNN (vs. $49.45\%$), $68.36\%$ on AT-rFGSM (vs. $26.10\%$), $65.37\%$ on AT-MaxMA (vs. $37.18\%$), and $36.82\%$ on PAD-SMA (vs. $8.25\%$).
Second, we introduce $\sigma$-binary attack, a novel adversarial method specifically designed for binary-constrained domains. This attack leverages a differentiable approximation of the Hamming distance, which is a key metric for binary vectors, to facilitate an effective gradient-based optimization. Furthermore, it incorporates dynamic confidence adjustment, enabling the optimization process to efficiently explore regions near the decision boundary and enhance the likelihood of identifying viable binary solutions. By seamlessly integrating Prioritized Binary Rounding, $\sigma$-binary maintains strict adherence to binary feature constraints while effectively achieving adversarial objectives.
Our experimental evaluations demonstrate the superior performance of $\sigma$-binary IV-C. Against the non-robust baseline (DNN), $\sigma$-binary achieves an optimal attack success rate of $100\%$, consistent with most other attacks except Mimicry. On AT-rFGSMk, it attains $99.45\%$, surpassing PGD-$\ell_{2}$ ($89.21\%$) and $\sigma$-zero ($85.95\%$). Against AT-MaxMA, $\sigma$-binary achieves $96.62\%$, outperforming $\sigma$-zero ($92.56\%$) and PGD-$\ell_{2}$ ($91.75\%$). For PAD-SMA, the most robust defense, $\sigma$-binary attains $94.56\%$, significantly exceeding PGD-$\ell_{2}$ ($79.90\%$) and $\sigma$-zero ($79.64\%$).
Further analysis of attack success rates across varying perturbation budgets confirms that $\sigma$-binary consistently outperforms all other attacks across all evaluated defenses (Figure 2), underscoring its efficacy in binary-constrained domains.
Finally, we apply the $\sigma$-binary attack to evaluate the robustness of the eight defenses on the Malscan dataset IV-C. The baseline DNN and defenses with adversary detectors, such as KDE, DLA, DNN+, and ICNN, exhibit significant vulnerability, with attack success rates exceeding $90\%$ for fewer than 10 feature modifications and reaching $100\%$ with a 20-feature perturbation budget. In contrast, adversarially trained defenses such as AT-rFGSMk and AT-MaxMA show improved robustness at lower budgets, but remain susceptible under unrestricted perturbations, with success rates of $99.45\%$ and $96.62\%$, respectively. PAD-SMA, considered the most robust defense, demonstrates strong resilience in prior evaluations [23], achieving an accuracy above $83.45\%$ (equivalent to an attack success rate below $16.55\%$) against 15 different gradient-based attacks without perturbation limits. However, our evaluation reveals vulnerabilities, with $\sigma$-binary achieving $36.34\%$ success with fewer than 10 modified features, $89.79\%$ at a 50-feature budget, and $94.56\%$ under unrestricted perturbations.
Furthermore, we evaluate defenses under both non-attack conditions and oblivious attacks to provide a comprehensive comparison. Our findings indicate that while PAD-SMA exhibits marginally greater robustness under high perturbation levels compared to the second most robust defense (AT-MaxMA), this comes at the cost of significant computational overhead and reduced accuracy for benign samples. These findings emphasize the trade-offs among various evaluation criteria and underscore the necessity to align defenses with specific threat models and operational requirements.
@@
I-B Paper outline.
The remainder of the paper is organized as follows: Section II provides a review of the preliminaries. Section III details the proposed methods, including Prioritized Binary Rounding and $\sigma$-binary. Section IV describes the experimental setup and presents results. Section V reviews related studies on evasion attacks and defenses. Finally, Section VI presents the key findings and outlines directions for future research.
II PRELIMINARIES
@@ II-A Notations
  • A feature extraction function $\phi:Z\to X$ maps samples from the problem space to a $d$-dimensional discrete feature space $X\subset\mathbb{R}^{d}$.
  • A malware detector $f:Z\to Y$ assigns a label $y\in Y=\{0,1\}$, where 00 represents benign and $1$ indicates malicious. The detector is defined as $f(z)=\varphi_{\theta}(\phi(z))$, where $\varphi_{\theta}:X\to Y$ is a ML model with parameters $\theta$.
  • An adversary detector $g:Z\to\mathbb{R}$ identifies adversarial examples. It operates in the feature space as $g(z)=\psi_{\vartheta}(\phi(z))$, where $\psi_{\vartheta}$ is a secondary model with learnable parameters $\vartheta$. A sample is flagged as adversarial if $g(z)>\tau$, where $\tau$ is a predefined threshold.
@@
II-B Integrated Malware and Adversary Detectors
To enhance the robustness of malware detection, a secondary adversary detector $g$ is integrated with the primary malware detector $f$. The models are defined as:
@@ II-B Integrated Malware and Adversary Detectors
$\text{predict}(z)=\begin{cases}f(z),&\text{if }g(z)\leq\tau,\\ 1,&\text{if }g(z)>\tau\text{ and }f(z)=1,\\ \text{further action},&\text{if }g(z)>\tau\text{ and }f(z)=0.\end{cases}$ (3)
@@ II-C Evasion Attacks
In the feature space, this translates to perturbations $x^{\prime}=x+\delta_{x}$ such that:
@@
The interval $[\underline{u},\overline{u}]$ defines the feature space boundaries, where $\underline{u}$ and $\overline{u}$ represent the lower and upper limits, respectively. These constraints ensure that perturbations remain within the valid feature range, maintaining the feasibility of adversarial examples.
@@
To address the disconnect between the feature space and the problem space, an approximate inverse mapping function, $\tilde{\phi}^{-1}$, is utilized[27]. This function maps perturbations from the feature space back to the problem space, ensuring practical feasibility.
II-D Oblivious vs. Adaptive Attacks

### v2

@@ Abstract
Machine learning-based malware detectors effectively identify malicious Android applications, but remain highly vulnerable to evasion attacks, where minor binary feature perturbations can evade detection. Despite ongoing progress, the lack of comprehensive evaluation frameworks limits a clear understanding of defense robustness. This study introduces two contributions for adversarial evaluation in binary-constrained malware detection. First, we propose Prioritized Binary Rounding, an efficient technique that converts continuous perturbations into binary features while preserving a high attack success rate and minimal modification budgets. Second, we present the $\sigma$-binary attack, a gradient-based method tailored for binary domains, capable of producing sparse yet highly effective perturbations. Experimental results in the MalScan dataset show that $\sigma$-binary achieves consistently superior and stable performance. It surpasses state-of-the-art binary-domain approaches against both robust and non-robust defenses. Advanced defenses such as KDE, DLA, DNN+, and ICNN, when evaluated under the proposed $\sigma$-binary attack, exhibit over 90% attack success with fewer than 10 modified features and reach 100% under a 20-feature budget. Even PAD-SMA, previously reported to retain 83% accuracy, is compromised by $\sigma$-binary with 36.3% success under 10-features and 94.6% with no feature-budget constraints. These results highlight the need for precise evaluation tools such as the proposed $\sigma$-binary attack to uncover latent weaknesses and guide the design of truly robust malware detectors.
@@ I Introduction
The surge in mobile device attacks has escalated into a critical global concern. In 2023 alone, such attacks increased by 50% compared to the previous year. According to Kaspersky, their systems blocked 33.8 million instances of malware, adware, and riskware targeting mobile devices, with Android remaining the primary focus of cybercriminals [kaspersky2024]. This significant increase in Android malware underscores the urgent need for advanced detection strategies to counter increasingly sophisticated threats.
@@
Machine learning (ML) has become a cornerstone of Android malware detection, significantly improving scalability and detection precision [gao2024comprehensive, liu2024unraveling, arp2014drebin, he2022msdroid, liu2022deep]. Among ML-based approaches, static analysis methods, such as Drebin [arp2014drebin] and MaMaDroid [mariconti2016mamadroid], have demonstrated notable efficiency. These methods analyze application features without requiring execution, enabling rapid and scalable security assessments. Consequently, static analysis has become a foundational component of modern malware detection pipelines.
@@
Despite these advancements, ML-based malware detectors remain inherently vulnerable to adversarial attacks. These attacks exploit weaknesses in detection models by introducing carefully crafted perturbations to evade detection. Such attacks, which undermine reliability, are broadly categorized into evasion and poisoning attacks. Evasion attacks manipulate test examples during the deployment phase to bypass detection [grosse2017adversarial, huang2018adversarial, li2023black, bostani2024evadedroid, hu2022generating], whereas poisoning attacks compromise training datasets to degrade detector performance [chen2018automated, bala2022droidenemy]. This study focuses on evasion attacks because of their immediate implications for operational systems.
@@
Evasion attacks can be further divided into problem-space and feature-space attacks. Problem-space attacks modify malware applications directly, such as injecting benign ”gadgets” or altering application components to obscure malicious intent [pierazzi2020intriguing, chen2019android, tang2024dapadv]. In contrast, feature-space attacks modify feature vectors extracted from malware while preserving the functionality of the original sample [huang2018adversarial, grosse2017adversarial]. This study focuses exclusively on feature-space attacks due to their controlled evaluation capabilities and practical relevance.
@@
Defending against adversarial attacks remains a significant challenge. Adversarial training, which incorporates adversarial samples into training datasets, has shown promise in improving robustness [grosse2017adversarial]. However, these methods often fail to generalize to novel attack strategies, impose high computational costs, and risk overfitting to specific attacks [carlini2019evaluating]. Other defense mechanisms such as ensemble methods [smutz2016tree, li2020adversarial, ficco2021malware], auxiliary models such as Variational Autoencoders (VAEs) [li2021robust], and hybrid frameworks such as PAD-SMA [li2023pad] offer varying degrees of effectiveness. However, these approaches frequently involve trade-offs, including increased complexity and reduced detection accuracy.
@@
Evaluating the adversarial robustness of malware defenses in binary spaces presents significant challenges. Existing approaches face two major limitations. First, gradient-based attacks originally designed for continuous feature spaces are frequently adapted to binary domains using naive binarization techniques, such as deterministic or randomized rounding [huang2018adversarial]. These adaptations substantially reduce attack effectiveness and fail to accurately capture the discrete nature of binary features. Furthermore, the absence of gradient-based attack formulations explicitly designed for binary spaces stems from the inherent difficulty of optimizing within a non-convex and non-differentiable constraint space. Second, many existing evaluation methodologies rely on generalized attack settings that employ low-iteration gradient-based attacks with fixed hyperparameters across all defense models [li2023pad, li2021robust, zhou2023malpurifier, li2020adversarial, li2021framework]. This lack of adaptive optimization prevents such evaluations from effectively exposing model-specific vulnerabilities and leads to an overestimation of robustness. Notably, previous studies have reported that black-box attacks, such as Mimicry, can sometimes outperform gradient-based approaches in binary domains [li2020adversarial, li2023pad] despite having no access to model parameters or architecture. This counterintuitive result arises because gradient-based methods, although privileged with full model knowledge, struggle to optimize effectively under discrete and non-differentiable binary constraints, leading to suboptimal perturbations and reduced attack success.
To address these limitations, this paper introduces a unified and systematic framework for assessing and enhancing model robustness in binary-constrained malware detection. The proposed framework establishes a principled, binary-compliant evaluation process capable of generating functionally valid adversarial examples and providing a reliable foundation for comparative robustness analysis.
In this context, our study makes the following key contributions:
  1. 1.
    Prioritized Binary Rounding (PBR): We develop a rounding mechanism that effectively transforms continuous perturbations into valid binary feature representations while preserving high attack success rates and minimizing modification budgets.
  2. 2.
    $\sigma$-binary Attack: We introduce a novel gradient-based adversarial attack specifically tailored for binary domains, capable of generating sparse yet highly effective adversarial examples through a differentiable optimization process.
  3. 3.
    Comprehensive Experimental Evaluation: Extensive experiments on the MalScan dataset demonstrate that $\sigma$-binary achieves consistently superior and stable performance compared to state-of-the-art approaches in the binary domain, against both robust and non-robust defense conditions.
Overall, this work establishes a principled and reproducible foundation for adversarial robustness evaluation in malware detection, bridging the gap between continuous-space optimization methods and discrete, real-world feature constraints.
@@
I-A Paper outline
The remainder of the paper is organized as follows: Section II reviews related studies on evasion attacks and defenses. Section III provides a review of the preliminaries. Section IV details the proposed methods, including Prioritized Binary Rounding and $\sigma$-binary. Section V describes the experimental setup and presents results. Section VI discusses the main findings, interprets the observed trade-offs, and highlights the limitations of the study. Section VII analyzes the internal and external threats to validity and explains the measures taken to mitigate them. Finally, Section VIII presents the key findings and outlines directions for future research.
II Related Work
II-A Evasion Attacks in Malware Detection
Problem-space attacks alter application components such as the manifest file, Dalvik bytecode, or inject benign artifacts to bypass detection. Android HIV [chen2019android] perturbs Dalvik bytecode to evade detectors while maintaining functionality. Pierazzi et al. [pierazzi2020intriguing] introduced the injection of benign bytecode ”gadgets” to fool detection models. Tang et al. [tang2024dapadv] expanded this with DapAdv, leveraging a hierarchical attention mechanism to optimize code slicing for adversarial repackaging. Advanced techniques include reinforcement learning-based attacks like HRAT [zhao2021structural], and hybrid approaches using generative adversarial networks (GANs) [goodfellow2020generative] with coevolution algorithms [li2023black] have further improved the effectiveness of problem-space attacks. Black-box methods such as EvadeDroid [bostani2024evadedroid] demonstrate success with minimal feature knowledge, relying on random perturbations in benign code.
Feature-space attacks focus on manipulating extracted feature vectors without altering the application itself. Gradient-based methods [huang2018adversarial, grosse2017adversarial] optimize perturbations to evade detection. Ensemble-based techniques such as those in [li2020adversarial] combine diverse manipulation strategies for enhanced efficacy. GAN-based approaches, such as E-MalGAN [li2019adversarial] and the method proposed by Shahpasand et al. [shahpasand2019adversarial], craft adversarial examples that evade both malware detectors and adversarial detection systems.
II-B Defenses Against Evasion Attacks
Adversarial training is a well-established defense strategy that enhances model resilience by incorporating adversarial examples into the training process. While it proves effective against the perturbations encountered during training [grosse2017adversarial], this approach faces limitations, including vulnerability to novel attack styles, high computational overhead, and a tendency to overfit to specific attack methods [carlini2019evaluating].
Ensemble methods aim to enhance robustness by combining multiple classifiers or detectors. Smutz and Stavrou [smutz2016tree] proposed mutual agreement analysis, flagging uncertain predictions based on classifier disagreement. Ficco [ficco2021malware] introduced a dynamic ensemble system combining generic and specialized detectors. However, applying ensemble strategies to deep learning models remains challenging due to difficulties in coordinating diverse model predictions.
Defenses with auxiliary models utilize separate frameworks to identify adversarial examples or to purify inputs. Li et al. [li2021robust] employed a Variational Autoencoder (VAE) to detect adversarial malware via reconstruction errors, while Li et al. [li2018hashtran] combined hash transformations with denoising autoencoders to mitigate perturbations. Despite initial success, these approaches often struggle against adaptive attacks.
Recently, hybrid approaches have gained prominence, combining multiple defense strategies to address the limitations of individual methods. For instance PAD-SMA [li2023pad] integrates a malware detection system with an adversary detector, both strengthened through adversarial training with a diverse range of attacks. While this approach enhances robustness and generalization, it also incurs significant computational overhead and exhibits lower accuracy on benign samples, limiting its practicality in real-world scenarios.
III Preliminaries
@@ III-A Notations
  • A feature extraction function $\phi:Z\to X$ maps samples from the problem space to a $d$-dimensional discrete feature-space $X\subset\mathbb{R}^{d}$.
  • A malware detector $f:Z\to Y$ assigns a label $y\in Y=\{0,1\}$, where 0 represents benign and $1$ indicates malicious. The detector is defined as $f(z)=\varphi_{\theta}(\phi(z))$, where $\varphi_{\theta}:X\to Y$ is a ML model with parameters $\theta$.
  • An adversary detector $g:Z\to\mathbb{R}$ identifies adversarial examples. It operates in the feature-space as $g(z)=\psi_{\vartheta}(\phi(z))$, where $\psi_{\vartheta}$ is a secondary model with learnable parameters $\vartheta$. A sample is flagged as adversarial if $g(z)>\tau$, where $\tau$ is a predefined threshold.
@@
III-B Integrated Malware and Adversary Detectors
To enhance the robustness of malware detection, a secondary adversary detector $g$ is integrated with the primary malware detector $f$. The models are defined as.
@@ III-B Integrated Malware and Adversary Detectors
[…]
@@ III-C Evasion Attacks
In the feature-space, this translates to perturbations $x^{\prime}=x+\delta_{x}$ such that:
@@
The interval $\underline{u},\overline{u}]$ defines the feature-space boundaries, where $\underline{u}$ and $\overline{u}$ represent the lower and upper limits, respectively. These constraints ensure that perturbations remain within the valid feature range, maintaining the feasibility of adversarial examples.
@@
To address the disconnect between the feature-space and the problem space, an approximate inverse mapping function, $\tilde{\phi}^{-1}$, is utilized [vsrndic2014practical]. This function maps perturbations from the feature-space back to the problem space, ensuring practical feasibility.
III-D Adversarial Attack Methods
We consider a comprehensive set of attacks spanning greedy, gradient-based, and ensemble strategies. These methods provide a diverse benchmark for assessing defense robustness in binary malware feature spaces.
Projected Gradient Descent (PGD).
PGD [madry2017towards] performs iterative gradient ascent under a constraint projection:
$\delta^{(t+1)}=\Pi_{\mathcal{C}}\big(\delta^{(t)}+\alpha\nabla_{\delta}F(\theta,x+\delta^{(t)},y)\big),$
where $\Pi_{\mathcal{C}}$ ensures feasibility, and the gradient is normalized under $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ norms (PGD-$\ell_{p}$). PGD is widely adopted for its strong, iterative optimization performance.
Fast Gradient Sign Method (FGSM).
FGSM [goodfellow2014explaining] performs a single-step update along the sign of the gradient:
$x^{\prime}=\operatorname{Proj}_{[\underline{u},\,\overline{u}]}\big(x+\varepsilon\cdot\operatorname{sign}(\nabla_{x}F(\theta,x,y))\big),$
where $\varepsilon$ controls perturbation magnitude. It serves as a fast, baseline adversarial method and as a precursor to iterative variants.
Randomized FGSM (rFGSM).
rFGSM [huang2018adversarial] extends FGSM by introducing a random initialization and applying randomized rounding for discrete inputs, increasing attack diversity and success in binary domains.
Bit Gradient Ascent (BGA) and Bit Coordinate Ascent (BCA).
Both BGA and BCA [grosse2017adversarial, huang2018adversarial] target binary features directly. BGA flips all bits with gradients exceeding a dynamic threshold, while BCA greedily flips the single bit with the largest positive gradient at each step. These methods are computationally efficient yet highly effective for discrete domains.
Carlini–Wagner (CW) Attack.
The CW attack [carlini2017towards] minimizes the $\ell_{2}$-norm of perturbations while enforcing misclassification through constrained optimization, producing minimal yet effective feature modifications.
Mixture and Ensemble Attacks.
Mixture-based attacks [li2020adversarial, li2023pad] combine multiple attack strategies to enhance strength and transferability. The Max Mixture of Attacks (MaxMA) selects the most adversarial example among multiple candidates; Iterative MaxMA (iMaxMA) repeats this process iteratively; and the Stepwise Mixture of Attacks (SMA) applies a staged combination of attacks with adaptive step sizes and norms.
Mimicry Attack.
Mimicry attacks [biggio2013evasion, vsrndic2014practical] are gradient-free strategies that perturb malware to resemble benign samples. By iteratively querying the target model and adjusting features, they achieve evasion without requiring access to model internals.
$\sigma$-zero Attack
The $\sigma$-zero attack [cina2024sigma] is designed to minimize the number of modified features by approximating the $\ell_{0}$-norm and applying adaptive projections. It is effective in generating highly sparse perturbations, challenging defenses that assume adversarial examples require extensive feature modifications.
III-E Oblivious vs. Adaptive Attacks